# Retail Central Bank Digital Currency: Motivations, Opportunities, and Mistakes

### v2

@@ 2.3 The Eurozone: the European Central Bank
    However, given that retail payments in the EU are increasingly online rather than offline, with the value transacted electronically now exceeding €240 trillion per year [64], it is fair to say that the need for privacy is precisely about payments in which at least one party is online, and the proposal by the EC suggests that for transactions with the digital euro that are not “fully-offline”, that is, in which neither party is connected to a network beyond each other, the degree of privacy will match the degree of privacy offered by prevailing electronic transactions facilitated by banks and credit providers, such as card payments. Unfortunately, the privacy offered by such institutions is “privacy by promise”, as those institutions have access to information specifying not only how consumers spend their money, but also metadata, such as location and device information, when they do. It is fair to say that consumers are subject to profiling and surveillance whenever they spend their money in the prevailing electronic transactions referenced by the EC proposal, so to say that consumers would have the same degree of privacy as they would have in such transactions is to say that consumers would not have privacy when they use the digital euro.
@@
    Certified hardware for “fully-offline” payments. A careful reading shows that the reference to “offline” payments in the proposal for the digital euro really refers only to “fully-offline” transactions wherein both parties are completely disconnected from any third parties. It has been theoretically established that fair exchange between two parties is impossible without the involvement of a trusted third party [65], so in practice, the assumption that “fully-offline” payments are possible would necessitate the involvement of a third-party actor at the point of sale. This is ostensibly achievable if at least one party to the transaction has a device that carries out the will of a third party in preference to the will of either of the transacting parties, although in practice that implies the use of so-called “trusted computing” (also known as a “secure element”) [66, 67] that can attest to a third party that the device had or had not taken a certain action. This kind of remote attestation has been roundly criticised by technical experts and others who advocate for the rights of users exercise control over their own devices [68]. We have previously argued that it would be unreasonable to impose certified hardware requirements on non-custodial wallets [55]:
@@ 3 Themes
As can be seen from the well-publicised digital currency proposals around the world, some common assumptions have begun to materialise. Whether these assumptions result from the influence of incumbent businesses, or the efforts of trans-national regulatory organisations, or the echo chamber resulting from the sharing of ideas by central banks, or other geopolitical factors, their similarities are no coincidence.
@@ 5 Accounts all the time
  • Geopolitical interests. Similarly, governments are keen to pursue data collection in the interest of cooperation with international law enforcement activities, and national governments also have an incentive to use international rule-making organisations to project power beyond their borders. Consider the recommendations of the Financial Action Task Force (FATF), and Recommendation 16 (the “Travel Rule”) in particular, which specifies that both the payer and recipient of electronic transactions must generally be discoverable [107]. A recommendation adopted in 1996 also states: “Countries should further encourage in general the development of modern and secure techniques of money management […] as a means to encourage the replacement of cash transfers” [108]. In 2022, the FATF published a recommendation explicitly stating that the Travel Rule would require financial institutions “to share relevant originator and beneficiary information alongside virtual asset transactions” and characterising “unhosted wallets” as an “emerging risk” [109].
@@ 6.1 Competition and Innovation
One central question is whether a particular direction for the design of digital currency could be expected to nurture greater competition and innovation than any other. Were the models for digital currency based upon custodial accounts to be adopted, then digital currency would be more akin to bank deposits than to cash, and it would be likely that regulated financial institutions, mostly consumer-facing banks, would be the only or primary entities to provide services to users, even when there is no technical reason for which this must be the case. This approach can be expected to protect the market structure for electronic payments of today, and we do not envisage any major differences. In contrast, were an approach to digital currency based upon tokens and non-custodial wallets to be adopted instead, then digital currency would be more akin to cash. For the purpose of retail payments, not only would banks, incumbent financial institutions, and money services businesses continue to provide essential services, but other kinds of businesses and organisations would have a place in the value chain as well, including those that provide ancillary services for the purposes of facilitating payments, providing hardware, maintaining software, offering better user experiences, and even offering custody for digital assets, perhaps coupled with security and management services. The diversity in the range and types of services, as well as the competition and innovation afforded by lower barriers to entry, would enhance the efficiency and value delivered to users and consumers. Through the lens of competition, it is clear that designs oriented around adapting cash for the digital age, are preferable to those that seek to build a new back-end for custodial business models. An approach based on tokens and non-custodial wallets would offer a greater diversity in the kinds and types of organisations that can service the interface between citizens and the digital economy. It would also support a way for people to choose between service providers on an as-needed basis, without being locked into account-based relationships that can be used as an instrument of surveillance and control, thus strengthening the potential for competition and thus innovation in servicing retail payments in the economy.
@@ 7 The future of money
We suggest that the core desiderata and design characteristics of public infrastructure should be agreed upon by consensus and should position the individual user at the centre, taking their interests into account as non-negotiable requirements and foundational constraints, rather than treating them as easily curtailed pawns that the G7 and others have so casually sets aside in furtherance of a narrow set of interests.
@@ Acknowledgements
The authors acknowledge the UCL Future of Money Initiative and the Systemic Risk Centre of the London School of Economics.

### v3

@@ 2.3 The Eurozone: the European Central Bank
    However, given that retail payments in the EU are increasingly online rather than offline, with the value transacted electronically now exceeding €240 trillion per year [64], it is fair to say that the need for privacy is precisely about payments in which at least one party is online, and the proposal by the EC suggests that for transactions with the digital euro that are not “fully-offline”, that is, in which neither party is connected to a network beyond each other, the degree of privacy will match the degree of privacy offered by prevailing electronic transactions facilitated by banks and credit providers, such as card payments. Unfortunately, the privacy offered by such institutions is actually a limited form of confidentiality, or privacy by promise, as those institutions have access to information specifying not only how consumers spend their money, but also metadata, such as location and device information, when they do. It is fair to say that consumers are subject to profiling and surveillance whenever they spend their money in the prevailing electronic transactions referenced by the EC proposal, so to say that consumers would have the same degree of privacy as they would have in such transactions is to say that consumers would not have privacy when they use the digital euro.
@@
    Certified hardware for “fully-offline” payments. A careful reading shows that the reference to “offline” payments in the proposal for the digital euro really refers only to “fully-offline” transactions wherein both parties are completely disconnected from any third parties. It has been theoretically established that fair exchange between two parties is impossible without the involvement of a trusted third party [65], so in practice, the assumption that “fully-offline” payments are possible would necessitate the involvement of a third-party actor at the point of sale. This is ostensibly achievable if at least one party to the transaction has a device that carries out the will of a third party in preference to the will of either of the transacting parties, although in practice that implies the use of so-called “trusted computing”: a “trusted execution environment” or “secure element” that can attest to a third party that the device had or had not taken a certain action [66, 67]. This kind of remote attestation has been routinely criticised by technical experts and others who advocate for the rights of users to exercise control over their own devices [68]. We have previously argued that it would be unreasonable to impose certified hardware requirements on non-custodial wallets [55]:
@@
Thematic element Digital dollar Digital pound Digital euro
Stated motivation improved efficiency and convenience for retail payments economic growth, digital transformation financial stability, monetary sovereignty, strategic autonomy
Use of accounts users access CBDC via accounts with intermediaries PIPs provide “wallets” that are de facto accounts service providers offer “digital euro payment accounts”
Privacy privacy is subordinate to a transparency requirement; payers can be de-anonymised payers can be de-anonymised no privacy for “online” payments; limited confidentiality for “offline” payments inside the security envelope of a service provider
Custody service providers hold funds in accounts wallets operated by PIPs hold funds service providers hold funds in accounts
Role of identity services used to access accounts would be “identity verified” an “alias service” links pseudonyms to a verified identity “digital euro payment accounts” are subject to identification requirements
Holding limits holding limits are applied by intermediaries on a per-user basis holding limits are applied to CBDC wallets holding limits are applied to digital euro payment acounts
Table 1: Common themes observed among the CBDC proposals.
@@ 3 Themes
As can be seen from the well-publicised digital currency proposals around the world (see Table 1), some common assumptions have begun to materialise, prompting us to seek an explanation. We have observed that incumbent businesses have joined CBDC research programmes orchestrated by central banks, thus seizing the opportunity to shape the narrative toward their existing business models. In addition, trans-national regulatory organisations, such as the Financial Action Task Force (FATF), have reacted unfavourably to non-custodial and privacy-enhancing electronic payment systems. Meanwhile, central banks have used forums, such as the BIS, to share ideas about the design of CBDC systems without much public debate, thus creating an echo chamber. Whether the common assumptions about the design of CBDC result from a combination of these or other geopolitical factors, their similarities are no coincidence.
@@ 5 Accounts all the time
  • Geopolitical interests. Similarly, governments are keen to pursue data collection in the interest of cooperation with international law enforcement activities, and national governments also have an incentive to use international rule-making organisations to project power beyond their borders. Consider the recommendations of the FATF, and Recommendation 16 (the “Travel Rule”) in particular, which specifies that both the payer and recipient of electronic transactions must generally be discoverable [107]. A recommendation adopted in 1996 also states: “Countries should further encourage in general the development of modern and secure techniques of money management […] as a means to encourage the replacement of cash transfers” [108]. In 2022, the FATF published a recommendation explicitly stating that the Travel Rule would require financial institutions “to share relevant originator and beneficiary information alongside virtual asset transactions” and characterising “unhosted wallets” as an “emerging risk” [109].
@@ 6.1 Competition and Innovation
One central question is whether a particular direction for the design of digital currency could be expected to nurture greater competition and innovation than any other. Were the models for digital currency based upon custodial accounts to be adopted, then digital currency would be more akin to bank deposits than to cash, and it would be likely that regulated financial institutions, mostly consumer-facing banks, would be the only or primary entities to provide services to users, even when there is no technical reason for which this must be the case. This approach can be expected to protect the market structure for electronic payments of today, and we do not envisage any major differences. In contrast, were an approach to digital currency based upon tokens and non-custodial wallets to be adopted instead, then digital currency would be more akin to cash. For the purpose of retail payments, not only would banks, incumbent financial institutions, and money services businesses continue to provide essential services, but other kinds of businesses and organisations would have a place in the value chain as well, including those that provide ancillary services for the purposes of facilitating payments, providing hardware, maintaining software, offering better user experiences, and even offering custody for digital assets, perhaps coupled with security and management services. The diversity in the range and types of services, as well as the competition and innovation afforded by lower barriers to entry, would enhance the efficiency and value delivered to users and consumers. Through the lens of competition, it is clear that designs oriented around adapting cash for the digital age are preferable to those that seek to build a new back-end for custodial business models. An approach based on tokens and non-custodial wallets would offer a greater diversity in the kinds and types of organisations that can service the interface between citizens and the digital economy. It would also support a way for people to choose between service providers on an as-needed basis, without being locked into account-based relationships that can be used as an instrument of surveillance and control, thus strengthening the potential for competition and thus innovation in servicing retail payments in the economy.
@@ 7 The future of money
Breaking through the noise requires building upon the research that exposes the unsoundness in the foundations underpinning the prevailing proposals for the design of CBDC. For example, Hyoung-kyu Chey observed the difference between a “state theory of money”, which posits that the foundational role of money is its value as a unit of account, and a “commodity theory of money”, which posits that the foundational role of money is its value as a medium of exchange [131]. While the prevailing CBDC proposals seem to assume that the primary role of money is to serve as a unit of account, experience with cryptocurrencies suggests that the need for CBDC arises instead from the role of money as a medium of exchange, highlighting a mismatch between the theoretical framework from which account-based CBDC proposals seem to arise and the constraints facing retail consumers as they pay for goods and services in the digital age. Similarly, although some of the justifications given for the proposed designs for CBDC reflect fears about the disintermediation of banks, research by Bibi and Canelli suggests that such fears are overwrought [132].
@@
We suggest that the core desiderata and design characteristics of public infrastructure should be agreed upon by consensus and should position the individual user at the centre, taking their interests into account as non-negotiable requirements and foundational constraints, rather than treating them as easily curtailed pawns that the G7 and others have so casually set aside in furtherance of a narrow set of interests.
@@ Acknowledgements
The authors acknowledge the support and encouragement of anonymous reviewers, anonymous central bankers, the UCL Future of Money Initiative, and the Systemic Risk Centre at the London School of Economics.